OpenStudy (anonymous):

2x + 1 5 or -3x > -9 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

sam (.sam.):

You missed an arrow in your question

OpenStudy (anonymous):

no it wwas the > , equal to sign , i cant do it on my laptop :o

sam (.sam.):

2x + 1>5 or -3x > -9

OpenStudy (anonymous):

but theres a line under it _> liike that i guess

sam (.sam.):

Solve for both \[2x+1\geq5 \\ \\ 2x \geq 4 \\ \\ x \geq 2\] ------------------------------------------- \[-3 \geq-9 \] Divide both sides by -3 flips the sign \[x \leq 3\]

OpenStudy (anonymous):

but how do you explain the graphing?

sam (.sam.):

Rewrite that as \[2 \leq x \leq 3\]

sam (.sam.):

x is in the region \[2 \leq x \leq 3\]

OpenStudy (anonymous):

but we have to graph it on a number line

sam (.sam.):



OpenStudy (anonymous):

thank you soooooooooo much !

sam (.sam.):

Welcome :)